OpenStudy (anonymous):

The height of an object above the ground at time t is given by s=v0t - (g/2)t^2 where v0 is the initial velocity and g is the acceleration due to gravity. a) At what height is the object initially? s=__________ ^^ i'm not quite sure how to find this... :( b) How long is the object in the air before it hits the ground? t=________ ^^again, not very sure.. how would you set this equation up? :/ c) When will the object reach its maximum height? t=_______ d) What is that maximum height? t=_______ **not really sure how to set up this whole thing :( Please explain? THankss :)

OpenStudy (agent0smith):

s=v0t - (g/2)t^2 when t=0, what will s be? plug in t=0

OpenStudy (anonymous):

s=v0t-(g/2)0^2 s=v0t-0 ?

OpenStudy (agent0smith):

Plug in t=0 but do it better :P

OpenStudy (anonymous):

oh oops i forgot one! s=v0(0)-0= 0 ?

OpenStudy (agent0smith):

Good :) s=v0t - (g/2)t^2 Factor out the t... s=t( v0t - (g/2)t ) reaches the ground means s=0 0=t( v0t - (g/2)t ) this would be far easier with equation editor too.

OpenStudy (agent0smith):

It's on the ground at t = 0 and... v0t - (g/2)t = 0 try solving for t.

OpenStudy (anonymous):

t-(g/2)t=v0t ? is that how to start it?

OpenStudy (anonymous):

this is part b right? and part a was s=0 ? :)

OpenStudy (agent0smith):

yes ...oops my mistake v0 - (g/2)t = 0 -(g/2)t=v0 t = -2v0/g

OpenStudy (anonymous):

ohh okay haha :P so that's as simplified as it can get right?

OpenStudy (agent0smith):

yeah. It's too hard to write that w/o equation editor, it'll do c) When will the object reach its maximum height? t=_______ It'll basically be the time above, divided by 2: t = -2v0/g/2 = -v0/g

OpenStudy (anonymous):

yeah for some reason, the equation and draw tools still aren't working :(

OpenStudy (agent0smith):

oh my the last is the worst to do w/o the equation editor d) What is that maximum height? plug in the above t into s=v0t - (g/2)t^2 s=v0(-v0/g) - (g/2)(-v0/g)^2 ew. You can simplify but i don't want to...

OpenStudy (anonymous):

hahaha okay.. so would it become s=(-v0^2/g) - (g/2)(-v0^2/g^2) ? is that right? :/ and if so, would it simplify any more?

OpenStudy (agent0smith):

Wow, good job putting up with it! Yes :) but you can cancel off one of the g's in the second term there

OpenStudy (agent0smith):

s=(-v0^2/g) - (1/2)(-v0^2/g)

OpenStudy (anonymous):

ohhh okay awesome!! :) hehe thanks!!! :) the tools really need to start working again soon!!

OpenStudy (agent0smith):

Yes they really do :/